# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ 5.4. Software failure
System Vulnerabilities: System vulnerabilities in UAVs refer to weaknesses in the software architecture, coding logic, or configuration settings that can be exploited by attackers or cause unintended behavior. These vulnerabilities may include unvalidated inputs, buffer overflows, and insecure memory handling, which expose the system to risks like memory corruption and code injection (Pradhan et al., 2015; Wang et al., 2023a; Ma et al., 2024). UAVs operating in sensitive environments such as surveillance or defense are especially at risk. For example, attackers may exploit unpatched firmware or default access credentials to hijack UAV functions or inject malicious commands (Vamvakas et al., 2019). Some studies have also highlighted vulnerabilities in channel access mechanisms, where channel access attacks (CAA) can degrade the freshness and reliability of information by disrupting data transmission through protocol exploitation (Wang et al., 2023a). Additionally, system-level software faults, including bugs, logic errors, insecure configurations, integration issues within dynamic cyber-physical environments, or soft errors from environmental factors like radiation, can compromise functionality and security, disrupting service delivery and threatening mission continuity (Ren et al., 2017; Pradhan et al., 2015; Ma et al., 2024; Venkatagiri et al., 2018a; Qu et al., 2022; Chang et al., 2024). Software aging is another critical concern, as prolonged operation may introduce subtle faults such as memory leaks or degraded performance, which can eventually lead to system failure or ineffective execution of key functions like task offloading (Watanabe and Machida, 2022). Furthermore, complex detection algorithms, such as those based on YOLOv5s, are vulnerable to configuration errors and mispredictions, which may result in high false alarm rates and reduced reliability of UAV-based detection systems (Zhong et al., 2021; Zuo et al., 2022; Zhou et al., 2023). Deep learning-based flight control systems may also encounter software faults, such as incorrect parameterization or model misbehavior in neural schedulers or DNNs, potentially leading to unsafe or inaccurate navigation decisions (Chen et al., 2022a). Monocular quadrotors also suffer from software limitations in depth estimation, where the absence of 3D information leads to unreliable obstacle avoidance, especially in unfamiliar or unstructured environments (Yang et al., 2019). Some UAV systems implement fault-tolerant mechanisms that enable rapid re-planning in response to actuator or sensor faults and allow seamless path switching through precomputed recovery scenarios in the event of unexpected software errors (Hirota et al., 2020a). An attacker could also intercept unencrypted telemetry data and manipulate it to alter flight paths or collect confidential information. Regular software audits, secure coding practices, and real-time intrusion detection systems are essential to mitigate these risks.
@@
System Crashes: UAV system crashes can occur due to software bugs, memory leaks, unexpected sensor inputs, or processing resources overload (Ma et al., 2024; Amin and Obermaisser, 2018a; Hirota et al., 2020a). These crashes may freeze the onboard computer, trigger automatic reboot, shut down critical flight functions or cause navigation-related faults (Kramer and Stubberud, 2021). For example, a software loop that does not terminate might consume all of the system memory, leading to a crash during flight. If recovery systems are not well implemented, the UAV may fall or crash uncontrollably. Edge-enabled UAVs are also vulnerable to system crashes if compute offloading mechanisms fail or if the drone is left without sufficient onboard fallback, potentially halting autonomous functions during critical missions (Liu et al., 2019). In multi-UAV edge orchestration frameworks, such as those used for collaborative video analytics, poorly optimized task scheduling or simultaneous high-load processing may overwhelm system resources, resulting in unresponsive drones or mid-mission system reboots (Qu et al., 2022). Misconfigured control logic at edge terminals, especially when not detected by testing, can result in severe flight instabilities or control system crashes, as demonstrated in fuzzing-based evaluations of UAV safety risks (Chang et al., 2024). Deep learning based detection modules like AIMED-Net, when deployed on resource-constrained edge platforms, may experience inference failures or overload under real-time constraints, potentially leading to unresponsive systems or control crashes during UAV operation. Comparative evaluations of onboard processors further highlight that under heavy computational loads, low-tier processors can become bottlenecks, increasing the risk of midmission system instability or crash (Pan et al., 2024; Jennings et al., 2022). Detection algorithms based on YOLOv5s can also suffer from configuration errors or inference instability, and in some cases, software crashes during execution, further increasing the risk of unresponsiveness in UAV detection systems(Shang et al., 2023; Taleb et al., 2023a; Zhang et al., 2021c; Grasso et al., 2022b; Ritchie et al., 2020; Lin et al., 2021b; Mo et al., 2022; Gupta et al., 2021; Hong et al., 2022). Other high-performance visual networks, such as VDTNet, may also introduce instability during edge-based detection and tracking tasks when deployed on resource-limited UAV platforms, potentially resulting in dropped detection, inference delays, or runtime crashes during intruder pursuit missions (Zhou et al., 2024d). System crashes can also stem from environmental hardware faults, energy depletion, unreliable communication links, or faults in AI-based detection and GPS/BDS-based attitude determination systems, which may cause incorrect orientation estimates and lead to UAV instability or crashes (Zhong et al., 2021; Nguyen et al., 2023; Zhang and Li, 2018; Kaymaz et al., 2024; Ahamad and Mishra, 2024). Additionally, inaccurate estimation of stability and control derivatives from simulated flight data can result in poorly tuned controllers, increasing the risk of instability or system crash during real-world UAV operations (Goel et al., 2021). A detailed fault tree of such failures in edge-based UAV orchestration has also been discussed in recent studies (Bai et al., 2023b). Reliability testing, hardware-in-the-loop simulations, and watchdog timers help manage the risk of software crashes.
@@ 5.5. Environmental impact
Adverse Weather Conditions: Environmental factors such as extremes of wind, rain, snow, fog, temperature, and dynamic air traffic can dramatically affect UAV performance and cause network disconnections that disrupt communication and data flow(Wang et al., 2021c, d; Gallego-Madrid et al., 2020; Hossain et al., 2017; Pradhan et al., 2015; Haber et al., 2021a). For example, strong winds can destabilize flight, while rain can interfere with electronics and sensors that are not adequately protected (Yang et al., 2022; Lee et al., 2022a; Sankarasrinivasan et al., 2015). Cold weather reduces battery efficiency, limiting flight time, and icy conditions may add weight or damage propellers. Environmental factors also affect thermal sensor accuracy (Barnawi et al., 2021). UAVs used for delivery or inspection must be ruggedized or equipped with weather-resilient sensors capable of adjusting flight behavior based on real-time input. Similarly, environmental factors such as sunlight glare, waves, and varying water color pose challenges for detection, while limited data availability further risks accuracy (Mcilwaine and Rivas Casado, 2021). Wind gusts may also cause sudden attitude changes, while electromagnetic interference, particularly near industrial zones or high voltage lines, can corrupt communication signals or disrupt GNSS-based navigation (Yang et al., 2022). Cloud cover and adverse conditions can cause data gaps or errors in mapping, reducing the reliability of timely agricultural monitoring (Fan et al., 2023). In this context, weather forecasting and adaptive flight planning systems are essential to ensure mission reliability and operational continuity. Additionally, UAVs operating at high altitudes are vulnerable to soft errors caused by radiation exposure, which can induce silent data corruptions in memory or logic circuits, affecting onboard computation and processes such as video summarization. These transient faults may go undetected, leading to reliability degradation during critical missions. To mitigate these risks, systems often require redundancy mechanisms, error-correcting codes, or radiation-hardened hardware components (Venkatagiri et al., 2018b). Studies have shown that lightweight deep learning models such as MobileNet strike an effective balance between inference performance and power consumption, making them suitable for real-time trail detection in UAVs operating in adverse conditions (Kumar and Jannesari, 2023).
@@
Terrain Challenges: Terrain features such as mountains, forests, urban environments, or bodies of water can obstruct signals or limit UAV maneuverability, directly affecting operational reliability (Dai et al., 2018). GPS signals may be reflected or blocked in canyons or dense urban areas, causing multipath errors and reduced navigational accuracy. Maintaining centimeter-level positioning accuracy becomes even more difficult under such conditions, particularly when using low-cost RTK-GNSS base stations (Rokaha et al., 2019). Tall buildings create radio control dead zones, while navigating tight spaces increases collision risks. In VANET-assisted UAV systems, urban obstacles such as buildings can disrupt radio waves, leading to frequent disconnections and failed message delivery (Paranjothi et al., 2019; Bakhtiari et al., 2020a). UAVs face coverage challenges due to occlusions in geometrically complex areas like urban or mountainous terrains, which can degrade image quality (Gao et al., 2022; Peng et al., 2023). Limited onboard energy further necessitates efficient path planning to maximize mission effectiveness (Dai et al., 2018). Additionally, detection systems deployed in remote or challenging terrains often face network disconnections, which can delay alerts and responses critical for low-altitude security operations (Kim et al., 2024b; Faraci et al., 2023a). Computational limits and channel obstacles hinder UAV task offloading, especially for dependent tasks modeled as directed acyclic graphs (Mao et al., 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments (Shang et al., 2023). In addition to signal-related issues, hardware faults caused by environmental stress or mechanical wear in rugged terrains may lead to individual UAV failures during missions (Lv et al., 2024). Furthermore, network disconnections in obstructed environments degrade coordination among UAVs, potentially disrupting formation maintenance and task execution (Wang et al., 2021f). Obstacles at ground level can also interrupt UAV-to-IoT communication, while limited battery capacity and hardware faults may disrupt edge computing services; laser-based power transfer has been proposed to stabilize such operations (Hadi and Ghazizadeh, 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments. Multi-UAV path planning also becomes difficult in uncertain, cluttered terrains where real-time obstacle avoidance and inter-UAV cooperation are critical for collision prevention and mission success (Zhou et al., 2024a).
@@ 5.6. Operational failures
Human error is a leading cause of UAV accidents, often arising during mission planning, parameter input (e.g., altitude, geofencing), incorrect mode selection, or delayed responses in emergencies (Xu and Zhai, 2020). Typical scenarios include unintentionally disabling obstacle avoidance or launching missions with insufficient battery charge, which can result in system loss, property damage, or injury. Regular training and procedural checklists help mitigate such errors (Hou et al., 2020b). Recent studies highlight a range of human-involved or human-impacting operational vulnerabilities. In MEC-enabled UAV systems, operator-induced mistakes or insufficient monitoring can degrade system performance under tight latency and energy constraints (Yu et al., 2021; Liu et al., 2023c; Yuan et al., 2022; Falcão et al., 2023a). Human misjudgment during UAV placement, route optimization, and offloading decisions may cause connectivity failures and task delays, especially in dynamic and uncertain network conditions (Kim et al., 2024b; Mao et al., 2023; Faraci et al., 2023b; Thapliyal et al., 2024a). Inaccurate assessment of battery status or communication strength also contributes to mission failure, particularly when UAVs serve dual roles (e.g., computing and relaying) (Zhang et al., 2024d; Xu and Zhai, 2020; Manzoor et al., 2021). Scenarios involving limited pilot observation further complicate UAV status awareness, increasing the likelihood of unsafe maneuvers or late interventions (Jeong et al., 2020a). In high-demand or post-disaster environments, human error in managing energy and computing resources may result in service outages or imbalanced workloads (He et al., 2023b; Hui et al., 2024; Falcão et al., 2023a; Li et al., 2024e; Kaleem et al., 2019b). Moreover, inadequate responses to user-level interference (e.g., overloading UAV bandwidth) or security threats (e.g., eavesdropping) may stem from policy gaps or oversight limitations (Ozger et al., 2018a; Rao and Naidu, 2024; Ouyang et al., 2023; Ju et al., 2024a; Tang et al., 2023; Amos et al., 2021; Shen et al., 2023a). Finally, system-level miscoordination due to delayed or misinformed human actions, such as poor handoff management or overutilized UAVs, leads to reduced QoE and compromised safety (Wang et al., 2020a; Vamvakas et al., 2019; Asim et al., 2022). These scenarios underscore the importance of designing robust, human-aware systems, which include real-time feedback, operator training, and semi-autonomous error prevention mechanisms.
@@
Maintenance Issues: Inadequate or skipped maintenance can lead to undetected issues such as loose connectors, worn-out rotors, degraded solder joints, or battery defects. Over time, these minor faults may escalate into critical in-flight failures, such as mid-air blade breakage. Ensuring dependability of UAV-based systems, therefore, demands regular inspections, maintenance documentation, and onboard diagnostic tools. Maintenance concerns in drone swarms involve node-level failures, communication, structural, or mission-related issues that degrade collective reliability; complex network modeling is used to evaluate their impact (Guo et al., 2022b). UAV inspection systems face hardware wear and safety hazards, especially when image resolution or network stability limits detection reliability (Jo et al., 2018; Wu et al., 2024). Environmental factors such as weather and mechanical degradation introduce hardware faults that disrupt coordination and task execution, particularly in dynamic terrains (Lv et al., 2024). Effective maintenance planning hinges on accurate failure forecasting. Time series models like Holt-Winters improve predictive accuracy, reducing downtime and safety risks (Yang et al., 2017). Engine and component malfunctions in UAVs or aircraft, coupled with disconnections in diagnostic links, lead to increased latency and unreliable maintenance actions (Ren et al., 2017; Zhang, 2020). Software bugs and vulnerabilities further necessitate continuous maintenance in cyber-physical and IIoT-based UAV systems (Pradhan et al., 2015; Lin et al., 2021b). System performance can also be affected by faults in RFID tags, wearable computing devices, or embedded processors, where network disconnections and energy limits hinder tracking and diagnostic accuracy (Zhang and Zhang, 2024; Patnayak et al., 2021a). In UAV-assisted robotic inspections, long-distance signal loss and energy constraints complicate real-time data processing and HD video transmission (Shams et al., 2023). For offshore wind farm monitoring, limited battery life and long-range travel demand optimized maintenance routes and inspection scheduling (Huang and Wang, 2024).

### v2

@@ 5.4. Software failure
System Vulnerabilities: System vulnerabilities in UAVs refer to weaknesses in the software architecture, coding logic, or configuration settings that can be exploited by attackers or cause unintended behavior. These vulnerabilities may include unvalidated inputs, buffer overflows, and insecure memory handling, which expose the system to risks like memory corruption and code injection (Pradhan et al., 2015; Wang et al., 2023a; Ma et al., 2024). UAVs operating in sensitive environments such as surveillance or defense are especially at risk. For example, attackers may exploit unpatched firmware or default access credentials to hijack UAV functions or inject malicious commands (Vamvakas et al., 2019). Some studies have also highlighted vulnerabilities in channel access mechanisms, where channel access attacks (CAA) can degrade the freshness and reliability of information by disrupting data transmission through protocol exploitation (Wang et al., 2023a). Additionally, system-level software faults, including bugs, logic errors, insecure configurations, integration issues within dynamic cyber-physical environments, or soft errors from environmental factors like radiation, can compromise functionality and security, disrupting service delivery and threatening mission continuity (Ren et al., 2017; Pradhan et al., 2015; Ma et al., 2024; Venkatagiri et al., 2018; Qu et al., 2022; Chang et al., 2024). Software aging is another critical concern, as prolonged operation may introduce subtle faults such as memory leaks or degraded performance, which can eventually lead to system failure or ineffective execution of key functions like task offloading (Watanabe and Machida, 2022). Furthermore, complex detection algorithms, such as those based on YOLOv5s, are vulnerable to configuration errors and mispredictions, which may result in high false alarm rates and reduced reliability of UAV-based detection systems (Zhong et al., 2021; Zuo et al., 2022; Zhou et al., 2023). Deep learning-based flight control systems may also encounter software faults, such as incorrect parameterization or model misbehavior in neural schedulers or DNNs, potentially leading to unsafe or inaccurate navigation decisions (Chen et al., 2022a). Monocular quadrotors also suffer from software limitations in depth estimation, where the absence of 3D information leads to unreliable obstacle avoidance, especially in unfamiliar or unstructured environments (Yang et al., 2019). Some UAV systems implement fault-tolerant mechanisms that enable rapid re-planning in response to actuator or sensor faults and allow seamless path switching through precomputed recovery scenarios in the event of unexpected software errors (Hirota et al., 2020). An attacker could also intercept unencrypted telemetry data and manipulate it to alter flight paths or collect confidential information. Regular software audits, secure coding practices, and real-time intrusion detection systems are essential to mitigate these risks.
@@
System Crashes: UAV system crashes can occur due to software bugs, memory leaks, unexpected sensor inputs, or processing resources overload (Ma et al., 2024; Amin and Obermaisser, 2018; Hirota et al., 2020). These crashes may freeze the onboard computer, trigger automatic reboot, shut down critical flight functions or cause navigation-related faults (Kramer and Stubberud, 2021). For example, a software loop that does not terminate might consume all of the system memory, leading to a crash during flight. If recovery systems are not well implemented, the UAV may fall or crash uncontrollably. Edge-enabled UAVs are also vulnerable to system crashes if compute offloading mechanisms fail or if the drone is left without sufficient onboard fallback, potentially halting autonomous functions during critical missions (Liu et al., 2019). In multi-UAV edge orchestration frameworks, such as those used for collaborative video analytics, poorly optimized task scheduling or simultaneous high-load processing may overwhelm system resources, resulting in unresponsive drones or mid-mission system reboots (Qu et al., 2022). Misconfigured control logic at edge terminals, especially when not detected by testing, can result in severe flight instabilities or control system crashes, as demonstrated in fuzzing-based evaluations of UAV safety risks (Chang et al., 2024). Deep learning based detection modules like AIMED-Net, when deployed on resource-constrained edge platforms, may experience inference failures or overload under real-time constraints, potentially leading to unresponsive systems or control crashes during UAV operation. Comparative evaluations of onboard processors further highlight that under heavy computational loads, low-tier processors can become bottlenecks, increasing the risk of midmission system instability or crash (Pan et al., 2024; Jennings et al., 2022). Detection algorithms based on YOLOv5s can also suffer from configuration errors or inference instability, and in some cases, software crashes during execution, further increasing the risk of unresponsiveness in UAV detection systems(Shang et al., 2023; Taleb et al., 2023; Zhang et al., 2021c; Grasso et al., 2022b; Ritchie et al., 2020; Lin et al., 2021; Mo et al., 2022; Gupta et al., 2021; Hong et al., 2022). Other high-performance visual networks, such as VDTNet, may also introduce instability during edge-based detection and tracking tasks when deployed on resource-limited UAV platforms, potentially resulting in dropped detection, inference delays, or runtime crashes during intruder pursuit missions (Zhou et al., 2024c). System crashes can also stem from environmental hardware faults, energy depletion, unreliable communication links, or faults in AI-based detection and GPS/BDS-based attitude determination systems, which may cause incorrect orientation estimates and lead to UAV instability or crashes (Zhong et al., 2021; Nguyen et al., 2023; Zhang and Li, 2018; Kaymaz et al., 2024; Ahamad and Mishra, 2024). Additionally, inaccurate estimation of stability and control derivatives from simulated flight data can result in poorly tuned controllers, increasing the risk of instability or system crash during real-world UAV operations (Goel et al., 2021). A detailed fault tree of such failures in edge-based UAV orchestration has also been discussed in recent studies (Bai et al., 2023a). Reliability testing, hardware-in-the-loop simulations, and watchdog timers help manage the risk of software crashes.
@@ 5.5. Environmental impact
Adverse Weather Conditions: Environmental factors such as extremes of wind, rain, snow, fog, temperature, and dynamic air traffic can dramatically affect UAV performance and cause network disconnections that disrupt communication and data flow(Wang et al., 2021b, b; Gallego-Madrid et al., 2020; Hossain et al., 2017; Pradhan et al., 2015; Haber et al., 2021). For example, strong winds can destabilize flight, while rain can interfere with electronics and sensors that are not adequately protected (Yang et al., 2022; Lee et al., 2022a; Sankarasrinivasan et al., 2015). Cold weather reduces battery efficiency, limiting flight time, and icy conditions may add weight or damage propellers. Environmental factors also affect thermal sensor accuracy (Barnawi et al., 2021). UAVs used for delivery or inspection must be ruggedized or equipped with weather-resilient sensors capable of adjusting flight behavior based on real-time input. Similarly, environmental factors such as sunlight glare, waves, and varying water color pose challenges for detection, while limited data availability further risks accuracy (Mcilwaine and Rivas Casado, 2021). Wind gusts may also cause sudden attitude changes, while electromagnetic interference, particularly near industrial zones or high voltage lines, can corrupt communication signals or disrupt GNSS-based navigation (Yang et al., 2022). Cloud cover and adverse conditions can cause data gaps or errors in mapping, reducing the reliability of timely agricultural monitoring (Fan et al., 2023). In this context, weather forecasting and adaptive flight planning systems are essential to ensure mission reliability and operational continuity. Additionally, UAVs operating at high altitudes are vulnerable to soft errors caused by radiation exposure, which can induce silent data corruptions in memory or logic circuits, affecting onboard computation and processes such as video summarization. These transient faults may go undetected, leading to reliability degradation during critical missions. To mitigate these risks, systems often require redundancy mechanisms, error-correcting codes, or radiation-hardened hardware components (Venkatagiri et al., 2018). Studies have shown that lightweight deep learning models such as MobileNet strike an effective balance between inference performance and power consumption, making them suitable for real-time trail detection in UAVs operating in adverse conditions (Kumar and Jannesari, 2023).
@@
Terrain Challenges: Terrain features such as mountains, forests, urban environments, or bodies of water can obstruct signals or limit UAV maneuverability, directly affecting operational reliability (Dai et al., 2018). GPS signals may be reflected or blocked in canyons or dense urban areas, causing multipath errors and reduced navigational accuracy. Maintaining centimeter-level positioning accuracy becomes even more difficult under such conditions, particularly when using low-cost RTK-GNSS base stations (Rokaha et al., 2019). Tall buildings create radio control dead zones, while navigating tight spaces increases collision risks. In VANET-assisted UAV systems, urban obstacles such as buildings can disrupt radio waves, leading to frequent disconnections and failed message delivery (Paranjothi et al., 2019; Bakhtiari et al., 2020). UAVs face coverage challenges due to occlusions in geometrically complex areas like urban or mountainous terrains, which can degrade image quality (Gao et al., 2022; Peng et al., 2023). Limited onboard energy further necessitates efficient path planning to maximize mission effectiveness (Dai et al., 2018). Additionally, detection systems deployed in remote or challenging terrains often face network disconnections, which can delay alerts and responses critical for low-altitude security operations (Kim et al., 2024b; Faraci et al., 2023). Computational limits and channel obstacles hinder UAV task offloading, especially for dependent tasks modeled as directed acyclic graphs (Mao et al., 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments (Shang et al., 2023). In addition to signal-related issues, hardware faults caused by environmental stress or mechanical wear in rugged terrains may lead to individual UAV failures during missions (Lv et al., 2024). Furthermore, network disconnections in obstructed environments degrade coordination among UAVs, potentially disrupting formation maintenance and task execution (Wang et al., 2021d). Obstacles at ground level can also interrupt UAV-to-IoT communication, while limited battery capacity and hardware faults may disrupt edge computing services; laser-based power transfer has been proposed to stabilize such operations (Hadi and Ghazizadeh, 2023). These challenges necessitate robust collision avoidance, terrain-following sensors like LiDAR, and autonomous systems to maintain reliable UAV operation in complex environments. Multi-UAV path planning also becomes difficult in uncertain, cluttered terrains where real-time obstacle avoidance and inter-UAV cooperation are critical for collision prevention and mission success (Zhou et al., 2024a).
@@ 5.6. Operational failures
Human error is a leading cause of UAV accidents, often arising during mission planning, parameter input (e.g., altitude, geofencing), incorrect mode selection, or delayed responses in emergencies (Xu and Zhai, 2020). Typical scenarios include unintentionally disabling obstacle avoidance or launching missions with insufficient battery charge, which can result in system loss, property damage, or injury. Regular training and procedural checklists help mitigate such errors (Hou et al., 2020a). Recent studies highlight a range of human-involved or human-impacting operational vulnerabilities. In MEC-enabled UAV systems, operator-induced mistakes or insufficient monitoring can degrade system performance under tight latency and energy constraints (Yu et al., 2021; Liu et al., 2023c; Yuan et al., 2022; Falcão et al., 2023). Human misjudgment during UAV placement, route optimization, and offloading decisions may cause connectivity failures and task delays, especially in dynamic and uncertain network conditions (Kim et al., 2024b; Mao et al., 2023; Faraci et al., 2023; Thapliyal et al., 2024). Inaccurate assessment of battery status or communication strength also contributes to mission failure, particularly when UAVs serve dual roles (e.g., computing and relaying) (Zhang et al., 2024d; Xu and Zhai, 2020; Manzoor et al., 2021). Scenarios involving limited pilot observation further complicate UAV status awareness, increasing the likelihood of unsafe maneuvers or late interventions (Jeong et al., 2020). In high-demand or post-disaster environments, human error in managing energy and computing resources may result in service outages or imbalanced workloads (He et al., 2023; Hui et al., 2024; Falcão et al., 2023; Li et al., 2024e; Kaleem et al., 2019). Moreover, inadequate responses to user-level interference (e.g., overloading UAV bandwidth) or security threats (e.g., eavesdropping) may stem from policy gaps or oversight limitations (Ozger et al., 2018; Rao and Naidu, 2024; Ouyang et al., 2023; Ju et al., 2024; Tang et al., 2023; Amos et al., 2021; Shen et al., 2023a). Finally, system-level miscoordination due to delayed or misinformed human actions, such as poor handoff management or overutilized UAVs, leads to reduced QoE and compromised safety (Wang et al., 2020a; Vamvakas et al., 2019; Asim et al., 2022). These scenarios underscore the importance of designing robust, human-aware systems, which include real-time feedback, operator training, and semi-autonomous error prevention mechanisms.
@@
Maintenance Issues: Inadequate or skipped maintenance can lead to undetected issues such as loose connectors, worn-out rotors, degraded solder joints, or battery defects. Over time, these minor faults may escalate into critical in-flight failures, such as mid-air blade breakage. Ensuring dependability of UAV-based systems, therefore, demands regular inspections, maintenance documentation, and onboard diagnostic tools. Maintenance concerns in drone swarms involve node-level failures, communication, structural, or mission-related issues that degrade collective reliability; complex network modeling is used to evaluate their impact (Guo et al., 2022b). UAV inspection systems face hardware wear and safety hazards, especially when image resolution or network stability limits detection reliability (Jo et al., 2018; Wu et al., 2024). Environmental factors such as weather and mechanical degradation introduce hardware faults that disrupt coordination and task execution, particularly in dynamic terrains (Lv et al., 2024). Effective maintenance planning hinges on accurate failure forecasting. Time series models like Holt-Winters improve predictive accuracy, reducing downtime and safety risks (Yang et al., 2017). Engine and component malfunctions in UAVs or aircraft, coupled with disconnections in diagnostic links, lead to increased latency and unreliable maintenance actions (Ren et al., 2017; Zhang, 2020). Software bugs and vulnerabilities further necessitate continuous maintenance in cyber-physical and IIoT-based UAV systems (Pradhan et al., 2015; Lin et al., 2021). System performance can also be affected by faults in RFID tags, wearable computing devices, or embedded processors, where network disconnections and energy limits hinder tracking and diagnostic accuracy (Zhang and Zhang, 2024; Patnayak et al., 2021). In UAV-assisted robotic inspections, long-distance signal loss and energy constraints complicate real-time data processing and HD video transmission (Shams et al., 2023). For offshore wind farm monitoring, limited battery life and long-range travel demand optimized maintenance routes and inspection scheduling (Huang and Wang, 2024).